OpenStudy (anonymous):

can somebody help me with this question? 2: let f be the function given by : f(x)=((x^4)-(16x^2))^(1/2) Find the slope of the line normal to the graph of f at x = 5

OpenStudy (anonymous):

anyone?

OpenStudy (mertsj):

\[y=x^4-4x\] \[y'=4x^3-4\]

OpenStudy (mertsj):

The slope of the tangent at x = 5 is 4(5)^3-4 = 496

OpenStudy (mertsj):

So the slope of the perpendicular is -1/496

OpenStudy (anonymous):

where did u get 4x from?

OpenStudy (mertsj):

\[\sqrt{16x ^{2}}=4x\]

OpenStudy (mertsj):

Or is this the problem? \[f(x)=\sqrt{x^4-16x^2}\]

OpenStudy (anonymous):

that was the problem and it was 16x^2

OpenStudy (mertsj):

So are you studying derivatives?

OpenStudy (anonymous):

yea

OpenStudy (mertsj):

Then take the derivative of the function.

OpenStudy (anonymous):

i did then do wat next?

OpenStudy (mertsj):

Replace x with 5 and evaluate.

OpenStudy (mertsj):

That will be the slope of the tangent. The slope of the normal will be the negative reciprocal of the slope of the tangent.

OpenStudy (anonymous):

i got 34/3 then the negative recip should be -(3/34) right?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

thxxx im finally done with my projext XD

OpenStudy (mertsj):

I got 34/45 for the slope of the tangent.

OpenStudy (anonymous):

let me redo it n c wat i did wrong

OpenStudy (anonymous):

i typed everything into the caculator n i got 340/30---> 34/3

OpenStudy (mertsj):

Your derivative must be wrong then. I put it into Wolf and got \[\frac{2x(x^2-8)}{\sqrt{x^2(x^2-16)}}\]

OpenStudy (anonymous):

it is i got (4x^3-32x)/ 2 sqaureroot x^4-16x^2

OpenStudy (anonymous):

\[f(x)= \sqrt{x^{4}}-\sqrt{16x ^{2}}\] this was the original problem

OpenStudy (mertsj):

So when I typed the problem and asked you if that was the problem and you said it was you really meant it wasn't?

OpenStudy (anonymous):

nvm ur answer was correct mines wrong

OpenStudy (anonymous):

wait so the slope is -(45/34) ?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

thx